OpenStudy (anonymous):

How to solve -50x^3 - 24x^2 + 2x

OpenStudy (aakashsudhakar):

When it equals zero? That would be... -50x^3 - 24x^2 + 2x = 0 25x^3 + 12x^2 - x = 0 25x^2 + 12x - 1 = 0 [x = 0] <-- One solution of x. x = [-b ± √((b^2) - 4(a)(c))]/2a a = 25; b = 12; c = -1 x = [-12 ± √(144 + 100)]/50 x = [-12 ± √(244)]/50 x = 0.07241; x = -0.55241 Total solutions of x: x = -0.55241 x = 0 x = 0.07241